Equipment purchased at the beginning of the fiscal year for $60,000 is expected to have a useful life of 5 years, or 1,000 operating hours, and a residual value of $10,000. Compute the depreciation for the first and second years of use by each of the following methods:

(a)

straight-line

Year 1

Year 2

(b)

units-of-production (200 hours first year; 5000 hours second year)

Year 1

Year 2

(c)

declining-balance at twice the straight-line rate

Year 1

Year 2
Answer #1

Answer

  • Requirement [a]

A

Cost

$            60,000.00

B

Residual Value

$            10,000.00

C=A - B

Depreciable base

$            50,000.00

D

Life [in years]

5

E=C/D

Annual SLM depreciation

$            10,000.00

Year 1 = $ 10,000
Year 2 = $ 10,000

  • Requirement [b]

A

Cost

$            60,000.00

B

Residual Value

$            10,000.00

C=A - B

Depreciable base

$            50,000.00

D

Usage

                        1,000

E = C/D

Depreciation per hour

$                    50.00

Year 1 = 200 hrs x $ 50 = $ 10,000
Year 2 = 500 hours x $ 50 = $ 25,000

  • Requirement [c]

A

Cost

$            60,000.00

B

Residual Value

$            10,000.00

C=A - B

Depreciable base

$            50,000.00

D

Life [in years]

5

E=C/D

Annual SLM depreciation

$            10,000.00

F=E/C

SLM Rate

20.00%

G=F x 2

DDB Rate

40.00%

Year

Beginning Book Value

Depreciation rate

Depreciation expense

Ending Book Value

1

$             60,000.00

40.00%

$         24,000.00

$             36,000.00

2

$              36,000.00

40.00%

$         14,400.00

$             21,600.00

Year 1 = $ 24,000
Year 2 = $ 14,400